OpenStudy (maddie82502):

What is the solution to the equation? 4(x – 3) = 7x + 24 A. x = –12 B. x = 12 C. x = –36 D. x = 4

OpenStudy (maddie82502):

@texaschic101

pooja195 (pooja195):

 Distribute

pooja195 (pooja195):

4 times x =? 4 times -3=?

OpenStudy (maddie82502):

4*1=4 and 4*3=12

pooja195 (pooja195):

Good so 4x-12=7x+24 Now add 12 to both sides 24+12=?

OpenStudy (maddie82502):

36

pooja195 (pooja195):

Good 4x=7x+36 4-7=?

OpenStudy (maddie82502):

3

pooja195 (pooja195):

you get -3 \[\huge~\rm~-3x=36\] Divide both sides by -3 x=36 divided by -3

OpenStudy (maddie82502):

1 and 12

pooja195 (pooja195):

good so x=-12